# Mapping the Cloud-Driven Atmospheric Dynamics & Chemistry of an Isolated Exoplanet Analog with Harmonic Signatures

### v1

@@ 4.1.1 NIRISS/SOSS Harmonic Results
The lowest-order periodicity across NIRISS/SOSS wavelengths is the harmonic at 2.4 h, consistent with SIMP J0136’s rotation (Yang et al., 2016). These harmonics tend to have the highest uncertainty as the period is of similar order as the observational period.
CO bands, and to a lesser extent H2O, show retrieved periods that consistently fall between the fundamental rotation (, 2.4 h) and its first harmonic (, 1.2 h), producing beating patterns with characteristic periodicities spanning 1.2 to 2.4 h at pressure levels ranging from to bar. These beating patterns were also seen in Apai et al. (2017, 2021b); Fuda et al. (2024) for WISE 1049 B and Zhou et al. (2022) for VHS 1256 b.
@@ 4.2 Vertical Variability Mapping via Harmonic Waves
Time-resolved, multi-spectral data such as those collected by JWST NIRISS and NIRSpec instruments provide the opportunity to map the dynamic atmospheric physics of brown dwarfs and exoplanets at varying pressure levels. Using a method first demonstrated in Akhmetshyn et al. (2025), we use the best-fit harmonic models (Figure 2, Middle Row), along with a model contribution function (see §3.4), to create vertical flux variability heat maps (pressure vs. time) of the atmosphere of SIMP J0136. These vertical variability maps demonstrate the flux variability at each atmospheric layer, the interaction between layers, and unique spectral contributions to the object’s dynamic atmosphere.
@@ 4.2.1 Composite Vertical Variability Maps
To build the vertical variability maps, each 0.05 m bin in the contribution function is scaled by the corresponding best-fit lightcurve model for each timestep. The total flux at each pressure level across all wavelengths is then integrated, resulting in a 1-D flux array in pressure-level space. At each timestep, the process is then repeated; followed by normalization of flux at each pressure level, resulting in a composite vertical variability map (Figure 5) in terms of % deviation.
@@
SIMP J0136’s composite vertical variability maps of the NIRISS/SOSS and NIRSpec/PRISM data reveal both similarities in the atmospheric structure observed at each epoch and significant changes between the two observations (separated by h or rotations). Both NIRISS/SOSS and NIRSpec/PRISM maps show two dynamic interacting atmospheric layers with time-varying boundaries: (1) a lower pressure (higher altitude) overlying layer comprised of lower order harmonics and (2) a deeper layer (bar) corresponding to expected cloud formation regions of forsterite and potentially iron (see e.g., Vos et al., 2023; McCarthy et al., 2024, 2025). The time-varying pressure level boundary between the deep and overlying atmospheric layers suggests variations in vertical cloud structure.
@@
However, significant atmospheric changes can be seen between the two observations, suggesting complex weather patterns evolving over a relatively short timescale. In the NIRISS/SOSS observation, the deep layer’s harmonic pattern is alternating single peaks of minima-maxima-minima followed by a double-peaked maximum. For the NIRSpec/PRISM observation, there are two pairs of double-peaked minima and maxima. The overlying layer evolves in complexity with the NIRSpec/PRISM observation, including a double-peaked maxima not observable in the NIRISS/SOSS data. Note that the NIRSpec/PRISM data extend to longer wavelengths and, therefore, lower pressure levels than the NIRISS/SOSS. As a result, the overlying layer in the NIRSpec/PRISM observation possesses more retrieved structure.
4.2.2 Spectrally-Unique Vertical Variability Maps
@@
The broad NIR spectral coverage of JWST NIRISS/SOSS and NIRSpec/PRISM allows us to map the unique variability of specific wavelengths tied to distinct sources of variability. For example, we can map the variability of wavelengths where iron ( to m) and forsterite ( to m) cloud modulation are expected to have the most effect (Vos et al., 2023; McCarthy et al., 2025) or observe variations in molecular absorption bands (H2O, CO, CH4).
@@
Spectrally-unique vertical variability maps showing variation in cloud modulation and molecular absorption can be seen in Figure 6. These maps are created by starting with the composite map (Figure 5) and then subtracting from it a second map comprising the wavelengths available for which we are not interested for a particular test, resulting in a spectrally-unique, % deviation map. This % deviation map is then scaled by the average S/N for the wavelengths of interest. To maximize the S/N, an additional round of harmonic fitting was performed with 8.71 min temporal binning (for both observations) and spectral binning based on the entire wavelength region of interest (e.g., 1.4 to 1.8m for forsterite clouds).
@@
The spectrally-unique variability maps for wavelengths corresponding to expected variations due to mineral cloud modulation (i.e., iron and forsterite, Vos et al., 2023) demonstrate clear harmonic patterns for both the NIRISS/SOSS and NIRSpec/PRISM datasets (see Figure 6, Top Row). The lower layer (bar) is suspected to be iron cloud modulation with maximum signal-to-noise (S/Nmax) deviations of 18.5 for NIRISS/SOSS and 16.1 for NIRSpec/PRISM observations. The upper forsterite layer (centered at 1 bar) is approximately anti-correlated with the underlying iron layer and has S/Nmax deviations of 16.3 and 26.5 for NIRISS/SOSS and NIRSpec/PRISM respectively.
@@
Variability with S/Nmax = 5.5 (NIRISS/SOSS) and 14.0 (NIRSpec/PRISM) in wavelengths corresponding to dominant H2O absorption bands ( to m, to m for NIRISS/SOSS and to m for NIRSpec/PRISM) is anti-correlated with the suspected to m forsterite cloud modulation for both data sets (see Figure 6, Second Row). H2O dark spots (enhanced absorption regions) anti-correlate with the to m dark spots while H2O bright spots (reduced absorption regions) correlate with the to m dark spots. H2O variability is strongest (highest S/N) at the bar pressure level, is constrained to mbar in the NIRSpec/PRISM data, but is relatively unconstrained in terms of a lower pressure bounds for NIRISS/SOSS data due to limited wavelength coverage.
Similar to H2O, variability in the CO absorption bands ( to m, with the NIRSpec/PRISM map also including the to m CO band) is anti-correlated with the to m modulation and shows strongest variability at the bar pressure level (Figure 6, Third Row). Evidence for variability is marginal (S/Nmax = 2.78) for the NIRISS/SOSS data but strong (S/Nmax = 14.0) for NIRSpec/PRISM observations. The NIRISS/SOSS map extends up to the mbar level with a wedge-like pattern. The NIRSpec/PRISM map, on the other hand, is constrained to bar pressure levels. This difference is again likely due to the constraining effect of the wider spectral coverage of NIRSpec/PRISM.
As opposed to the H2O and CO maps, the wavelengths corresponding to lower-altitude (deeper pressure levels) CH4 absorption bands ( to m, to m, to m) are correlated with to m modulation in both observations (see Figure 6, Fourth Row). Evidence for lower-altitude CH4 variation is weaker than for H2O and CO, with marginal evidence (S/Nmax = 2.9) for NIRISS/SOSS data and tentative evidence (S/Nmax = 3.6) for the NIRSpec/PRISM maps. As with the CO maps, the NIRISS/SOSS variability extends up to the mbar level while the NIRSpec/PRISM variability is constrained to the bar pressure level.
As the NIRSpec/PRISM data extends to 5.3m, it provides access to higher-altitude spectral features (as shown in the contribution function in Figures 3 and 4). The wavelengths from m bracket a CH4 feature reaching mbar pressure levels. The high-altitude CH4 map (see Figure 6, Fifth Row) has a distinct (and potentially anti-correlated) variability pattern (S/Nmax = 14.9) compared to the lower-altitude CH4 absorption bands. Longer wavelengths also include the relatively high-altitude spectral continuum (3.05 to 3.15 m, 3.6 to 3.8 m, see Figure 6, Bottom Row) with variability (S/Nmax = 6.0) appearing to be correlated with the to m modulation and anti-correlated with high-altitude CH4 variability.
@@ 5.1 Odd Harmonics in the Forsterite Cloud Layer
As initially identified by Akhmetshyn et al. (2025) for the NIRISS/SOSS dataset, odd harmonics may indicate a North/South hemispheric asymmetry in the surface map of SIMP J0136. Such asymmetry would also support a inclination (Cowan et al., 2013), as edge-on geometries would make such asymmetry less likely, helping to further constrain the current estimate of (Vos et al., 2017). Furthermore, this asymmetry will help lift hemispheric ambiguity in future Doppler imaging efforts with ELT-class telescopes with high-resolution spectrographs (see e.g., Plummer & Wang, 2023), making SIMP J0136 an ever more compelling target for such work.
Because brown dwarfs have been found to be redder, and therefore cloudier, at equatorial latitudes than polar ones (Vos et al., 2017; Suárez & Metchev, 2022; Suárez et al., 2023), the odd harmonics at the deepest pressure levels (corresponding to the cloud formation region) suggest the North/South asymmetry may lie in the mineral cloud modulation region and be concentrated near the equator. Multi-rotational NIR spectroscopic observations of SIMP J0136 would help to constrain higher order odd harmonics, decrease the uncertainty in retrieved harmonic parameters, and elucidate the dynamical structure of L/T transition objects.
@@ 5.2 Cloud-Driven Temperature and Chemical Structure
The anti-correlation between the 1.4 to 1.8m forsterite cloud modulation region and H2O and CO absorption bands, as well as the correlation of CH4 with the same cloud region (see Figure 6), suggests that forsterite clouds influence the atmospheric chemical structure of SIMP J0136, potentially by driving variations in temperature structure and chemical mixing ratios. H2O and CO2 display reduced absorption at and above the cloud formation level while CH4 exhibits enhanced absorption. The differing behavior of CH4 implies that the explanation for the anti-correlation of H2O and CO2 variability with the cloud layers cannot be completely explained due to clouds muting the observability of these absorption features.
We suggest two potential causes for the relationship between the variability in the cloud modulation region and molecular absorption bands: (1) Cloud modulation influences the temperature structure, and therefore, the chemical conditions of the atmosphere and (2) Cloud formation and dissipation controls the oxygen abundances at co-altitude and overlying pressure levels. Both mechanisms could play a part in determining the dynamic and chemical structure of the upper atmosphere of SIMP J0136.
Forsterite cloud modulation can be seen to drive the upper atmosphere temperature structure as shown by the correlation between the dark cloud feature in Figure 6 (Top Row, Right Column) and the high-altitude continuum in Figure 6 (Bottom Row, Right Column). Regions above the clouds are relatively cooler while gaps in the cloud layer allow heat to escape initially in vertical columns (see e.g., Showman et al., 2019; Tan & Showman, 2021a, b). As CO is the preferred carbon reservoir at higher temperatures (and contrastingly CH4 is the preferred reservoir at lower temperatures), cooler regions should exhibit depleted CO and enhanced CH4 absorption, matching our findings (Figure 6, Third and Fourth Rows) and contributing to carbon disequilibrium chemistry in SIMP J0136.
To understand the H2O variation in this context, we should consider the following equation (e.g., Visscher & Moses, 2011),
@@
The sequestration of oxygen in the forsterite cloud formation process may also factor into the observed anti-correlation of the cloud layer and CO/H2O. The chemistry governing the forsterite cloud formation process can be understood by the net thermochemical reaction (e.g., Visscher et al., 2010),
where and respectively denote a solid or liquid phase. The process consumes H2O to form forsterite, potentially explaining the anti-correlation seen in Figure 6 (Second Row). Depleted oxygen abundances at and above the cloud layer also favor CH4 over CO as the preferred carbon reservoir (see Equation 5), again explaining the observed carbon disequilibrium chemistry in SIMP J0136 as well as the chemical variation patterns in Figure 6 (Third and Fourth Rows).

### v3

@@ 4.1.1 NIRISS/SOSS Harmonic Results
The lowest-order periodicity across NIRISS/SOSS wavelengths is the harmonic at 2.4 h, consistent with SIMP J0136’s rotation (Yang2016). These harmonics tend to have the highest uncertainty as the period is of similar order as the observational period.
CO bands, and to a lesser extent H2O, show retrieved periods that consistently fall between the fundamental rotation (, 2.4 h) and its first harmonic (, 1.2 h), producing beating patterns with characteristic periodicities spanning 1.2 to 2.4 h at pressure levels ranging from to bar. These beating patterns were also seen in Apai2017; apai21; Fuda:2024 for WISE 1049 B and Zhou2022 for VHS 1256 b.
@@ 4.2 Vertical Variability Mapping via Harmonic Waves
Time-resolved, multi-spectral data such as those collected by JWST NIRISS and NIRSpec instruments provide the opportunity to map the dynamic atmospheric physics of brown dwarfs and exoplanets at varying pressure levels. Using a method first demonstrated in Akhmetshyn2025, we use the best-fit harmonic models (Figure 2, Middle Row), along with a model contribution function (see §3.4), to create vertical flux variability heat maps (pressure vs. time) of the atmosphere of SIMP J0136. These vertical variability maps demonstrate the flux variability at each atmospheric layer, the interaction between layers, and unique spectral contributions to the object’s dynamic atmosphere.
@@ 4.2.1 Composite Vertical Variability Maps
To build composite vertical variability maps, we start by scaling the model contribution function, , see Equation 5, by the normalized best-fit lightcurve model, , see Equation 1, for each spectral bin, , and time-step. At each pressure level, , we then numerically sum the scaled spectral contribution across wavelength bins to compute the pressure- and time- dependent variability,
@@
resulting in a 1-D flux array in pressure-level space. The computation is then repeated at each time-step, followed by normalization of flux at each pressure level, resulting in a composite vertical variability map (Figure 5, Top Row) in terms of % deviation. The composite deviation maps broadly match the dynamic spectra (variability maps) in Figure 2, and the vertical variability maps can be seen as a coordinate transformation of the flux heat map from time vs. wavelength to time vs. pressure.
To quantify the significance of the observed features in the retrieved maps, we scale the composite variability map (Figure 5, Top Row) by the S/N in pressure vs. time space (Figure 5, Top Row). The S/N map is computed in a similar manner as the composite map, using Equation 6, but the normalized lightcurve flux is replaced with the S/N for each spectral bin and time-step. Comparing the S/N values in Figure 5, Bottom Row to the binned values in Figure 1, Bottom Row, shows agreement for both NIRISS/SOSS and NIRSpec/PRISM datasets. The resulting S/N-weighted map (Figure 5, Bottom Row) shows the highest deviation () in the deeper pressure levels (bar) due to the high S/N in these regions. The shallow layer (bar) exhibits smaller due to its lower S/N despite possessing larger deviation.
@@
SIMP J0136’s composite vertical variability maps of the NIRISS/SOSS and NIRSpec/PRISM data reveal both similarities in the atmospheric structure observed at each epoch and apparent qualitative differences between the two observations (separated by h or rotations). Both NIRISS/SOSS and NIRSpec/PRISM maps show two dynamic interacting atmospheric layers with time-varying boundaries: (1) a lower pressure (higher altitude) overlying layer comprised of lower order harmonics and (2) a deeper layer (bar) corresponding to expected cloud formation regions of forsterite and potentially iron (see e.g., Vos2023; McCarthy:2024; McCarthy2025). The time-varying pressure level boundary between the deep and overlying atmospheric layers suggests variations in vertical cloud structure.
@@
However, visual differences between the two observations’ maps can be seen, suggesting complex weather patterns evolving over a relatively short timescale. In the NIRISS/SOSS observation, the deep layer’s harmonic pattern is alternating single peaks of minima-maxima-minima followed by a double-peaked maximum. For the NIRSpec/PRISM observation, there are two pairs of double-peaked minima and maxima. The overlying layer evolves in complexity with the NIRSpec/PRISM observation, including a double-peaked maxima not observable in the NIRISS/SOSS data. Note that the NIRSpec/PRISM data extend to longer wavelengths and, therefore, lower pressure levels than the NIRISS/SOSS. As a result, the overlying layer in the NIRSpec/PRISM observation possesses more retrieved structure.
4.2.2 Spectral Features’ Deviations from Bulk Variability
@@
The broad NIR spectral coverage of JWST NIRISS/SOSS and NIRSpec/PRISM allows us to map spectral features’ deviations from bulk (or composite) variability and investigate that deviation’s ties to distinct atmospheric phenomena. For example, we can determine how wavelengths corresponding to iron ( to m) and forsterite ( to m) cloud modulation (Vos2023; McCarthy2025) or known molecular absorption bands (H2O, CO, CH4) deviate from bulk variability over the period of observation.
@@
Maps depicting specific spectral regions’ deviations from bulk variability can be seen in Figure 6. These maps are created by starting with the composite map (Figure 5) and then subtracting from it a second map comprising the wavelengths available for which we are not interested for a particular test, resulting in a spectrally-unique, % deviation map. This % deviation map is then scaled by the average S/N for the wavelengths of interest. To maximize the S/N, an additional round of harmonic fitting was performed with 8.71 min temporal binning (for both observations) and spectral binning based on the entire wavelength region of interest (e.g., 1.4 to 1.8m for forsterite clouds).
It is important to note that maps showing deviations of individual spectral features from the bulk variability do not necessarily indicate bona fide, feature-specific variability. Rather, they reflect differences in the flux relative to the composite (bulk) signal. For example, a molecular absorption band may exhibit no intrinsic time-dependent behavior, but if bulk variability is subtracted from the spectral band light curves, the residuals would exhibit an induced anti-correlation. Nevertheless, these deviations from the bulk variability are indeed meaningful, as they highlight wavelengths with distinct morphologies from the overall spectrum.
@@
The deviation maps for wavelengths expected to correspond to mineral cloud modulation (i.e., iron and forsterite, Vos2023) demonstrate clear harmonic patterns for both the NIRISS/SOSS and NIRSpec/PRISM datasets (see Figure 6, Top Row). The lower layer (bar) is suspected to be iron cloud modulation with maximum signal-to-noise (S/Nmax) deviations of 18.5 for NIRISS/SOSS and 16.1 for NIRSpec/PRISM observations. The upper forsterite layer (centered at 1 bar) is approximately anti-correlated with the underlying iron layer and has S/Nmax deviations of 16.3 and 26.5 for NIRISS/SOSS and NIRSpec/PRISM respectively.
@@
Deviation with S/Nmax = 5.5 (NIRISS/SOSS) and 14.0 (NIRSpec/PRISM) in wavelengths corresponding to dominant H2O absorption bands ( to m, to m for NIRISS/SOSS and to m for NIRSpec/PRISM) is anti-correlated with the suspected to m forsterite cloud modulation for both data sets (see Figure 6, Second Row). H2O dark spots (enhanced absorption regions) anti-correlate with the to m dark spots while H2O bright spots (reduced absorption regions) correlate with the to m dark spots. H2O deviations are strongest (highest S/N) at the bar pressure level, and are constrained to mbar in the NIRSpec/PRISM data, but are relatively unconstrained in terms of a lower pressure bounds for NIRISS/SOSS data due to limited wavelength coverage.
Similar to H2O, deviations in the CO absorption bands ( to m, with the NIRSpec/PRISM map also including the to m CO band) are anti-correlated with the to m modulation and show the strongest deviations at the bar pressure level (Figure 6, Third Row). Evidence for deviation is marginal (S/Nmax = 2.78) for the NIRISS/SOSS data but strong (S/Nmax = 14.0) for NIRSpec/PRISM observations. The NIRISS/SOSS map extends up to the mbar level with a wedge-like pattern. The NIRSpec/PRISM map, on the other hand, is constrained to bar pressure levels. This difference is again likely due to the constraining effect of the wider spectral coverage of NIRSpec/PRISM.
As opposed to the H2O and CO maps, the wavelengths corresponding to lower-altitude (deeper pressure levels) CH4 absorption bands ( to m, to m, to m) are correlated with to m modulation in both observations (see Figure 6, Fourth Row). Evidence for lower-altitude CH4 deviation is weaker than for H2O and CO, with marginal evidence (S/Nmax = 2.9) for NIRISS/SOSS data and tentative evidence (S/Nmax = 3.6) for the NIRSpec/PRISM maps. As with the CO maps, the NIRISS/SOSS deviations extend up to the mbar level while the NIRSpec/PRISM deviations are constrained to the bar pressure level.
As the NIRSpec/PRISM data extends to 5.3m, it provides access to higher-altitude spectral features (as shown in the contribution function in Figures 3 and 4). The wavelengths from m bracket a CH4 feature reaching mbar pressure levels. The high-altitude CH4 map (see Figure 6, Fifth Row) has a distinct (and potentially anti-correlated) deviation pattern (S/Nmax = 14.9) compared to the lower-altitude CH4 absorption bands. Longer wavelengths also include the relatively high-altitude spectral continuum (3.05 to 3.15 m, 3.6 to 3.8 m, see Figure 6, Bottom Row) with deviations (S/Nmax = 6.0) appearing to be correlated with the to m modulation and anti-correlated with high-altitude CH4 deviations.
@@ 5.1 Odd Harmonics in the Forsterite Cloud Layer
As initially reported by Akhmetshyn2025, odd () harmonics in the NIRISS/SOSS data may indicate North/South asymmetry in the surface brightness map of SIMP J0136 (depending on the limb darkening law; luger21d). In the absence of limb darkening, even spherical-harmonic degrees () paired with odd azimuthal orders (which correspond to the longitudinal wavenumber ) signify a North–South asymmetry in the surface-brightness distribution (Cowan&Fuentes2013), an assumption adopted by Akhmetshyn2025. The situation changes, however, once limb darkening is included: linear and quadratic limb–darkening laws effectively increase the contributing spherical-harmonic degree by one and two, respectively (luger21d). Consequently, a linear limb–darkening law can generate apparent odd- (odd-) signals even for a North–South symmetric map, whereas under a quadratic law—or in the limit of negligible limb darkening—the presence of such odd modes once again implies genuine hemispheric asymmetry. Additional constraints on the limb–darkening characteristics of SIMP J0136 and similar objects will therefore be essential for robust interpretation of their light-curve morphologies.
If the odd harmonics are indeed due to North/South asymmetry, then they also support an inclined viewing geometry (Cowan&Fuentes2013), consistent with the estimate of (Vos2017). Furthermore, this asymmetry could help lift hemispheric ambiguity in future Doppler imaging efforts with ELT-class telescopes with high-resolution spectrographs (see e.g., Plummer&Wang2023), making SIMP J0136 an ever more compelling target for such work.
Because brown dwarfs have been found to be redder, and therefore cloudier, at equatorial latitudes than polar ones (Vos2017; Suarez&Metchev2022; Suarez2023), the odd harmonics at the deepest pressure levels (corresponding to the cloud formation region) potentially suggest the North/South asymmetry may lie in the mineral cloud modulation region and be concentrated near the equator. Multi-rotational NIR spectroscopic observations of SIMP J0136 would help to constrain higher order odd harmonics, decrease the uncertainty in retrieved harmonic parameters, and elucidate the dynamical structure of L/T transition objects.
@@ 5.2 Cloud-Driven Temperature and Chemical Structure
The anti-correlation between the 1.4 to 1.8m forsterite cloud modulation region and H2O and CO absorption bands, as well as the correlation of CH4 with the same cloud region (see Figure 6), suggests that forsterite clouds influence the atmospheric chemical structure of SIMP J0136, potentially by driving variations in temperature structure and chemical mixing ratios. H2O and CO2 display reduced absorption at and above the cloud formation level while CH4 exhibits enhanced absorption. The differing behavior of CH4 implies that the explanation for the anti-correlation of H2O and CO2 deviations with the cloud layers cannot be completely explained due to clouds muting the observability of these absorption features.
We suggest two potential causes for the relationship between the deviations in the cloud modulation region and molecular absorption bands: (1) Cloud modulation influences the temperature structure, and therefore, the chemical conditions of the atmosphere and (2) Cloud formation and dissipation controls the oxygen abundances at co-altitude and overlying pressure levels. Both mechanisms could play a part in determining the dynamic and chemical structure of the upper atmosphere of SIMP J0136.
Forsterite cloud modulation can be seen to drive the upper atmosphere temperature structure as shown by the correlation between the dark cloud feature in Figure 6 (Top Row, Right Column) and the high-altitude continuum in Figure 6 (Bottom Row, Right Column). Regions above the clouds are relatively cooler while gaps in the cloud layer allow heat to escape initially in vertical columns (see e.g., Showman&Tan2019; tan21a; tan21b). As CO is the preferred carbon reservoir at higher temperatures (and contrastingly CH4 is the preferred reservoir at lower temperatures), cooler regions should exhibit depleted CO and enhanced CH4 absorption, matching our findings (Figure 6, Third and Fourth Rows) and contributing to carbon disequilibrium chemistry in SIMP J0136.
To understand the H2O deviations in this context, we should consider the following equation (e.g., Visscher2011),
@@
The sequestration of oxygen in the forsterite cloud formation process may also factor into the observed anti-correlation of the cloud layer and CO/H2O. The chemistry governing the forsterite cloud formation process can be understood by the net thermochemical reaction (e.g., Visscher2010),
where and respectively denote a solid or liquid phase. The process consumes H2O to form forsterite, potentially explaining the anti-correlation seen in Figure 6 (Second Row). Depleted oxygen abundances at and above the cloud layer also favor CH4 over CO as the preferred carbon reservoir (see Equation 7), again explaining the observed carbon disequilibrium chemistry in SIMP J0136 as well as the chemical deviation patterns in Figure 6 (Third and Fourth Rows).